# Clustering-based Low-Rank Matrix Approximation: An Adaptive Theoretical Analysis with Application to Data Compression

### v1

@@ VI Discussion and Limitations
To compare the performance of cluster-based and global SVD across different modalities, we computed the relative error within the ROI and outside it, along with the SSIM error and IoU error. In additionally, the EPI error was computed specifically within the tumor region to evaluate the preservation of structural details in the area of clinical relevance. Before computing these metrics, all reconstructed and ground truth images were resized to the same dimension. Furthermore, the compression factor was kept approximately the same for all modalities to allow for fair assessment of reconstruction quality. As shown in Table I, cluster-based SVD consistently outperformed global SVD across all medical image modalities. The X-ray image modality preserved and extracted important feature the most, especially at a higher compression factor, due to their homogeneous backgrounds and sharp edges. The consistently low errors in X-ray image modality suggest that cluster-based SVD is more suited for this modality. CT scan image modality benefited notably, with better preservation of subtle gradients important for diagnostic accuracy. In MRI, the presence of complex anatomical textures and high-frequency tumor regions complicates the clustering process, leading to moderate gain. For ultrasound images, both cluster-based and global SVD methods performed relatively poorly, with high relative, SSIM, and IoU errors. This is due to ultrasound’s low contrast and speckle noise which reduces patch similarity.
TABLE I: Comparison between cluster‐based and global SVD methods across different medical imaging modalities (MRI, ultrasound, CT scan, and X‑ray) and two randomly selected compression factors (6 and 40). For each modality and compression factor, the table reports the relative reconstruction error (Rel. Err), SSIM error, IoU error, and EPI error (all in %) within the ROI and the existing metrics outside the ROI.
@@
VII Conclusion and future work
@@
In this work, we introduced a cluster-based SVD framework to address the challenges of compressing large-scale data matrices while preserving and extracting critical structural details. Our study comprehensively compares global and cluster-based SVD. These methods were applied to medical images from different modalities, such as MRI, CT scan, ultrasound, and X-ray, to evaluate their effectiveness in preserving and extracting important features. Our analysis showed that the uniform compression enforced by global SVD fails to adapt to the local structural complexity of medical images. This often leads to visible artifacts and the loss of diagnostically important features, especially in heterogeneous regions such as tumors or lesions. In contrast, the cluster-based SVD method exploits local and non-local self-similarity to apply adaptive compression. This approach consistently improves image fidelity across modalities, enabling better preservation of fine structural details, as shown by higher PSNR, SSIM, IoU, and EPI values. However, our experimental results also highlight that the performance of cluster-based SVD is modality-dependent. Cluster-based SVD performed best in X-ray image modality, where the homogeneous background and sharp edges allowed for effective grouping and compression. On the other hand, it performed poorly on ultrasound image modality, where the presence of speckle noise and low contrast significantly reduced patch similarity, reducing the ability of cluster-based SVD to exploit localized low-rank structures effectively. Additionally, we provided a theoretical framework to quantify the compression factor. We showed how the overall compression factor is influenced by parameters such as patch size, $K$ and $\alpha$. Moreover, we demonstrated that there is not always a general global rank $r_{g}$ for which the cluster-based and global SVD compression factors are approximately the same. While cluster-based SVD has higher computational complexity than global SVD, the study demonstrated that these costs could be mitigated through parallel processing, making the approach feasible for large-scale data matrices.
@@
Future work will investigate alternative clustering methods that are less computationally expensive than k-means. While k-means is effective in grouping structurally similar patches, its iterative nature and sensitivity to initialization can lead to increased processing time, especially in high-resolution data matrices. Exploring faster clustering techniques may help reduce the overall computational burden while maintaining or improving clustering quality. Exploring clustering techniques, such as mini-batch k-means or hierarchical clustering with early stopping, may offer speed-ups without significantly compromising performance. Such improvements could make the proposed method more scalable, especially for high-resolution 3D medical data matrices and real-time clinical applications.
ACKNOWLEDGMENT
@@
VII-A Experiment 1: MRI
This section presents a detailed performance comparison between global SVD and cluster‐based SVD for an MRI image in Fig. 5(i). Fig. 9 compares edge preservation in global and cluster-based SVD reconstructions. Fig. 9(a) presents the original medical image, while Fig. 9(b) displays the edge map extracted from it, highlighting the natural structural details. Fig. 9(c) shows the edge map obtained from the image reconstructed using global SVD, showing that many fine edges have been lost or blurred. In contrast, Fig. 9(d) presents the edge map from the cluster-based SVD reconstruction, where edges are sharper and more consistent with the original image. The edge loss maps in Figs. 9(e) and (f) further emphasize this difference. Fig. 9(e) shows widespread edge loss throughout the image, whereas Fig. 9(f) demonstrates more localized and less severe loss, indicating better preservation of structural details.
@@ VII-A Experiment 1: MRI
VII-B Experiment 2: Ultrasound
VII-B1 Dataset Description
@@ VII-B2 Performance Evaluation
VII-C Experiment 3: CT scan
VII-C1 Dataset Description
@@ VII-C2 Performance Evaluation
VII-D Experiment 4: X-ray
VII-D1 Dataset Description

### v2

@@ 6 Discussion and Limitations
To compare the performance of cluster-based and global SVD across different modalities, we computed the relative error within the ROI and outside it, along with the SSIM error and IoU error. In additionally, the EPI error was computed specifically within the tumor region to evaluate the preservation of structural details in the area of clinical relevance. Before computing these metrics, all reconstructed and ground truth images were resized to the same dimension. Furthermore, the compression factor was kept approximately the same for all modalities to allow for fair assessment of reconstruction quality. As shown in Table 1, cluster-based SVD consistently outperformed global SVD across all medical image modalities. The X-ray image modality preserved and extracted important feature the most, especially at a higher compression factor, due to their homogeneous backgrounds and sharp edges. The consistently low errors in X-ray image modality suggest that cluster-based SVD is more suited for this modality. CT scan image modality benefited notably, with better preservation of subtle gradients important for diagnostic accuracy. In MRI, the presence of complex anatomical textures and high-frequency tumor regions complicates the clustering process, leading to moderate gain. For ultrasound images, both cluster-based and global SVD methods performed relatively poorly, with high relative, SSIM, and IoU errors. This is due to ultrasound’s low contrast and speckle noise which reduces patch similarity.
Table 1: Comparison between cluster‐based and global SVD methods across different medical imaging modalities (MRI, ultrasound, CT scan, and X‑ray) and two randomly selected compression factors (6 and 40). For each modality and compression factor, the table reports the relative reconstruction error (Rel. Err), SSIM error, IoU error, and EPI error (all in %) within the ROI and the existing metrics outside the ROI.
@@
7 Conclusion and future work
@@
In this work, we introduced a cluster-based SVD framework to address the challenges of compressing large-scale medical images while preserving and extracting critical structural details. Our study comprehensively compares global and cluster-based SVD. These methods were applied to medical images from different modalities, such as MRI, CT scan, ultrasound, and X-ray, to evaluate their effectiveness in preserving and extracting important features. Our analysis showed that the uniform compression enforced by global SVD fails to adapt to the local structural complexity of medical images. This often leads to visible artifacts and the loss of diagnostically important features, especially in heterogeneous regions such as tumors or lesions. In contrast, the cluster-based SVD method exploits local and non-local self-similarity to apply adaptive compression. This approach consistently improves image fidelity across modalities, enabling better preservation of fine structural details, as shown by higher PSNR, SSIM, IoU, and EPI values. However, our experimental results also highlight that the performance of cluster-based SVD is modality-dependent. Cluster-based SVD performed best in X-ray image modality, where the homogeneous background and sharp edges allowed for effective grouping and compression. On the other hand, it performed poorly on ultrasound image modality, where the presence of speckle noise and low contrast significantly reduced patch similarity, reducing the ability of cluster-based SVD to exploit localized low-rank structures effectively. Additionally, we provided a theoretical framework to quantify the compression factor. We showed how the overall compression factor is influenced by parameters such as patch size, $K$ and $\alpha$. Moreover, we demonstrated that there is not always a general global rank $r_{g}$ for which the cluster-based and global SVD compression factors are approximately the same. While cluster-based SVD has higher computational complexity than global SVD, the study demonstrated that these costs could be mitigated through parallel processing, making the approach feasible for large-scale medical images.
@@
Future work will investigate alternative clustering methods that are less computationally expensive than k-means. While k-means is effective in grouping structurally similar patches, its iterative nature and sensitivity to initialization can lead to increased processing time, especially in high-resolution medical images. Exploring faster clustering techniques may help reduce the overall computational burden while maintaining or improving clustering quality. Exploring clustering techniques, such as mini-batch k-means or hierarchical clustering with early stopping, may offer speed-ups without significantly compromising performance. Such improvements could make the proposed method more scalable, especially for high-resolution 3D medical images and real-time clinical applications.
[…]
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
@@ Appendix A Appendix
A.1 Experiment 1: MRI
This section presents a detailed performance comparison between global SVD and cluster‐based SVD for an MRI image in Fig. 5(i) of the main document. Fig. 9 compares edge preservation in global and cluster-based SVD reconstructions. Fig. 9(a) presents the original medical image, while Fig. 9(b) displays the edge map extracted from it, highlighting the natural structural details. Fig. 9(c) shows the edge map obtained from the image reconstructed using global SVD, showing that many fine edges have been lost or blurred. In contrast, Fig. 9(d) presents the edge map from the cluster-based SVD reconstruction, where edges are sharper and more consistent with the original image. The edge loss maps in Figs. 9(e) and (f) further emphasize this difference. Fig. 9(e) shows widespread edge loss throughout the image, whereas Fig. 9(f) demonstrates more localized and less severe loss, indicating better preservation of structural details.
@@ A.1 Experiment 1: MRI
A.2 Experiment 2: Ultrasound
A.2.1 Dataset Description
@@ A.2.2 Performance Evaluation
A.3 Experiment 3: CT scan
A.3.1 Dataset Description
@@ A.3.2 Performance Evaluation
A.4 Experiment 4: X-ray
A.4.1 Dataset Description